### v2

# Characterization of Deletion/Substitution Channel Capacity for Small Deletion and Substitution Probabilities

Abstract
We consider binary input deletion/substitution channels, which model certain channels with synchronization errors encountered in practice. Specifically, we focus on the regime of small deletion and substitution probabilities, and by extending an approach developed for the deletion-only channel, we obtain an asymptotic characterization of the channel capacity for independent and identically distributed deletion/substitution channels. We first present an upper bound on the capacity for arbitrary but fixed numbers of deletions and substitutions, and then we extend the result to the case of random deletions and substitutions. Our final result is as follows: The i.i.d. deletion/substitution channel capacity is approximately $1-H(p_{d})-H(p_{s})$, for $p_{d},p_{s}\approx 0$, where $p_{d}$ is the deletion probability, and $p_{s}$ is the substitution probability.
I Introduction
Channels with insertions, deletions, and substitutions are crucial in modeling certain communication systems with synchronization errors. For instance, when there is a clock mismatch in a wireless communication system, such errors are observed. In addition, bit-patterned media recording channels and DNA data storage channels exhibit such errors. In this paper, our focus is on the special case of channels experiencing bit losses (deletions) and bit errors (substitutions). Specifically, we are interested in independent, identically distributed (i.i.d.) deletions and substitutions at the regime of small deletion and substitution probabilities. Our objective is to characterize the channel capacity for this asymptotic regime.
Dobrushin established the information stability of channels with memoryless synchronization errors in [1], which is extended in [2] to synchronization errors governed by a Markov chain. However, obtaining the exact capacity has proved very difficult. Even for a channel with only i.i.d. deletions, the capacity is still unknown. As a result, the literature has focused on obtaining bounds on the capacity. Gallager’s work on sequential decoding for channels with synchronization errors provides the first lower bound on the capacity of a deletion channel [3], which states that for deletion probabilities of less than $1/2$, i.e., $p_{d}<1/2$, the capacity of i.i.d. binary deletion channel (BDC) is not smaller than that of a binary symmetric channel with a cross-over probability of $p_{d}$; that is,
$C_{d}\geq 1-H(p_{d}),$ (1)
where $H(\cdot)$ denotes the binary entropy function. Indeed, Gallager’s work is more general and establishes a lower bound on the capacity of i.i.d. insertion, deletion, and substitution error channels. For the case of deletion/substitution channels (i.e., with insertion probability of zero), the lower bound becomes:
$C_{d,s}\geq 1+p_{d}\log p_{d}+p_{s}\log p_{s}+p_{c}\log p_{c},$ (2)
where $p_{s}$ is the substitution probability, and $p_{c}=1-p_{d}-p_{s}$ is the probability of correct transmission. In the regime of asymptotically small synchronization errors, this can be lower bounded as
$C_{d,s}\geq 1-H(p_{d})-H(p_{s}).$ (3)
We will prove that this bound is asymptotically tight by establishing an upper bound on the capacity whose dominant terms match it in the asymptotic regime of $p_{d},p_{s}\rightarrow 0$. There are also other attempts to obtain lower bounds on the capacity of channels with synchronization errors. The authors in [4] employ tree codes to obtain lower bounds for insertion/deletion channels. In [5], a lower bound on the BDC capacity is obtained using the jigsaw-puzzle decoding approach, which is further improved in [6].
Only a limited number of upper bounds have been presented for the capacity of channels with synchronization errors. In [7], the authors provide two upper bounds on the BDC capacity: a numerically evaluated one for any arbitrary deletion probability and a simple closed-form one. In [8], numerical upper bounds are provided for deletion/substitution channels by exploiting an auxiliary system where the receiver is provided with side information. The authors extend their results to a non-binary deletion channel in [9], and to more general synchronization error channels with insertions, deletions, and substitutions in [10]. The authors of [11] improve the BDC lower bound in [6] as well as the upper bound in [8]. In [12], by converting the capacity problem into maximizing a univariate, real-valued, and often concave function, the authors obtain upper bounds for a general class of repetition channels.
Several works have focused on approximating the capacity of synchronization error channels for the asymptotic regime of small synchronization errors. In [13, 14], the BDC capacity is approximated using series expansions up to the second and third leading terms, respectively, by calculating the dominant terms of the capacity expansion for small deletion probabilities. Similar results are obtained for duplication channels [15] and insertion channels [16].
More in line with our work, the authors of [17] show that, for small deletion probabilities, the capacity can be upper-bounded as
$C\leq 1-(1-o(1))H(p_{d}).$ (4)
To obtain this bound, they first obtain lower and upper bounds on the success probability of a guesser that, given the received erroneous sequence, outputs the transmitted sequence along with the deletion pattern. To this end, assuming that the decoding is successful with a certain probability, they obtain bounds on the number of deletion patterns in a certain vicinity of the desired one that can be mistaken for it at the receiver. Finally, combining (4) with Gallager’s lower bound, they show that the BDC capacity approaches $1-H(p_{d})$ as $p_{d}\rightarrow 0$. The authors extend the result to the $k$-copy BDC in [18].
In this paper, we adapt the general methodology in [17] and extend it to channels with substitution errors as well as deletions. We first provide an upper bound on the capacity of deletion/substitution channels. To this end, we first fix the numbers of deletions and substitutions and assume that the receiver knows these numbers as side information (it knows the number of deletions anyway due to the received sequence length), and obtain an upper bound on the capacity, which is also an upper bound on the capacity of the case without side information since side information cannot decrease capacity. Combining with the result in [3], we obtain a tight approximation of the capacity when the deletion and substitution probabilities are asymptotically small. Finally, we extend the result to the case of random numbers of deletions and substitutions, and obtain a tight approximation for the capacity in the asymptotic regime of small deletion and substitution probabilities.
[…]
II System Model
We consider a deletion/substitution channel, where each bit is independently deleted with probability $p_{d}$, flipped (substituted) with probability $p_{s}$, or correctly transmitted with probability $1-p_{d}-p_{s}$. We consider a codebook $C\subseteq\{0,1\}^{n}$ consisting of messages of $n$ bits and of size $N=|C|$. We may think of a transmission pattern $A_{t}=a^{t}_{1},a^{t}_{2},...,a^{t}_{n-q_{d}}$ as a strictly increasing subsequence of $[n]:=\{1,2,...,n\}$, representing which bits are transmitted (not deleted), where $q_{d}$ is the number of deletions. Similarly, we denote a substitution pattern $A_{s}$ as a strictly increasing subsequence of $[n-q_{d}]$, $A_{s}=a^{s}_{1},a^{s}_{2},...,a^{s}_{q_{s}}$, where $q_{s}$ is the number of substitutions.
A deletion/substitution channel with deletion probability $p_{d}$ and substitution probability $p_{s}$ can be decomposed as a concatenation of a deletion channel with deletion probability $p_{d}$ followed by a substitution channel with substitution probability $p_{s}^{\prime}=\frac{p_{s}}{1-p_{d}}$, as depicted in Fig. 1, which we found to be easier to work with for our analytical derivations.
Figure 1: Decomposition of a deletion/substitution channel with deletion probability $p_{d}$ and substitution probability $p_{s}$ as a concatenation of a deletion channel with deletion probability $p_{d}$ followed by a substitution channel with substitution probability $p_{s}^{\prime}=\frac{p_{s}}{1-p_{d}}$.
Each realization of the deletion/insertion channel can be represented as a pattern pair $A=(A_{t},A_{s})$. For a string $X\in\{0,1\}^{n}$, $X_{A}$ represents the transmission of $X$ through a deletion/substitution channel with transmission and substitution patterns $A_{t}$ and $A_{s}$, respectively, in the obvious way: the bits of $X$ are deleted according to the transmission pattern $A_{t}$, then the remaining bits are flipped with the substitution pattern $A_{s}$.
III Fixing the numbers of deletions and substitutions
We aim to obtain an upper bound on the size of a uniquely decodable codebook, whose logarithm gives us an upper bound on the capacity. We first obtain the upper bound on the codebook size for fixed numbers of deletions and substitutions, assuming that the receiver knows them as side information. Note that the receiver is always aware of the number of deletions, as it is equal to the length loss of the received sequence, but not the number of substitutions.
To obtain the upper bound on the codebook size, we first obtain upper and lower bounds on the success probability of a guesser that, given a received sequence, guesses the transmitted codeword as well as the transmission (deletion) and substitution patterns. Then, we compare these upper and lower bounds to obtain an upper bound on the codebook size.
In the following, we provide an upper bound on the success probability of a general guesser with side information.
Lemma 1 ([17])
Let $\rho$ be a joint distribution over $S\times T$, for finite sets $S$ and $T$, such that the marginal distribution over $S$ is uniform. Let $g:T\rightarrow S$ be a function. Then, for $(a,b)$ following the joint distribution $\rho$, $Pr[g(b)=a]\leq\frac{|T|}{|S|}$.
Corollary 1
Using Lemma 1, for fixed numbers of deletions and substitutions, we obtain the following upper bound on the success probability of a general guesser:
$Pr_{Z\in U^{\mathcal{C}}}[g(Z_{A})=(Z,A)]\leq\frac{2^{n-q_{d}}}{N{n\choose q_{d}}{n-q_{d}\choose q_{s}}},$ (5)
since the possible number of transmitted codewords $Z$ is equal to the codebook size $N$ and the number of possible deletion/substitution patterns is (${n\choose q_{d}}{n-q_{d}\choose q_{s}}$. Also, since the output length is equal to $n-q_{d}$, the number of possible output sequences is upper bounded by $2^{n-q_{d}}$.
To obtain a lower bound on the success probability, we first need to obtain an upper bound on the number of neighboring patterns that can be mistaken for the actual one. To this end, we first upper bound the probability of two transmission/substitution pattern pairs leading to the same output for all possible inputs in the following lemma.
Lemma 2
Take any two transmission/substitution pattern pairs, $A=(A_{t},A_{s})$ and $B=(B_{t},B_{s})$. For uniformly random $X\in\{0,1\}^{n}$, $Pr_{X\in U^{\{0,1\}^{n}}}[X_{A}=X_{B}|(A_{t},A_{s}),(B_{t},B_{s})]\leq 2^{-|\Delta_{t}|}$ if $\delta_{s}\subseteq\Delta_{t}$, and zero otherwise, where $\Delta_{t}$ denotes the discrepancy set between the two transmission patterns and $\delta_{s}$ is the symmetric difference of the two substitution patterns [^1] Note that $\Delta_{t}$ is defined over the indices of the elements of the two sequences, not on their values.; i.e., $\Delta_{t}=\Delta(A_{t},B_{t})$ and $\delta_{s}=\delta(A_{s},B_{s})$.
Proof:
In [17], it is shown that in the case of only deletions, the probability of channel outputs being the same is upper bounded as $Pr_{X\in U^{\{0,1\}^{n}}}[X_{A_{t}}=X_{B_{t}}]\leq 2^{-|\Delta_{t}|}$. In the case of only substitutions, two outputs can be the same only if the two substitution patterns are exactly the same. However, that is not the case for the deletion/substitution scenario. In that case, as long as the substitution pattern difference set is a subset of the discrepancy set of the transmission patterns, the two outputs can be the same with the same probability as the two transmission patterns that lead to the same output. So, we have
$\begin{split}Pr_{X\in U^{\{0,1\}^{n}}}&[X_{A}=X_{B}|(A_{t},A_{s}),(B_{t},B_{s})]\\ &\begin{cases}=1,&\text{if }A=B\\ =0,&\text{if }A_{t}=B_{t},A_{s}\neq B_{s}\\ \leq 2^{-|\Delta_{t}|},&\text{if }A_{t}\neq B_{t},A_{s}=B_{s}\\ =0,&\text{if }A_{t}\neq B_{t},A_{s}\neq B_{s},\delta_{s}\not\subseteq\Delta_{t}\\ \leq 2^{-|\Delta_{t}|}.&\text{if }A_{t}\neq B_{t},A_{s}\neq B_{s},\delta_{s}\subseteq\Delta_{t}\\ \end{cases}\end{split}$ (6)
Simplifying this gives us the final expression in the lemma. ∎
Corollary 2
The expression in Lemma 1 can be further upper bounded as $Pr_{X\in U^{\{}0,1\}^{n}}[X_{A}=X_{B}|(A_{t},A_{s}),(B_{t},B_{s})]\leq 2^{-|\Delta_{t}|}$.
In the next lemma, we upper bound the number of patterns in a certain neighborhood of a given pattern that can be mistaken for it.
Lemma 3
For any pattern $A=(A_{t},A_{s})$ and integers $s$ and $t$, the number of patterns $B=(B_{t},B_{s})$ with $|\Delta_{t}|\leq t$ and $|\delta_{s}|\leq s$ that can be mistaken with $A$ is upper bounded by
$(s+1)(t+1){q_{s}+s\choose q_{s}}{t+s\choose s}{{q_{d}+t}\choose{q_{d}}}{{2q_{d}+t+1}\choose{2q_{d}+1}}.$ (7)
Proof:
From Lemma 1, two patterns can only be mistaken with each other if the set of their substitution differences is a subset of their transmission pattern differences; i.e., $\delta_{s}\subseteq\Delta_{t}$. The number of such patterns, $I$, can be expressed as
$\begin{split}&I=|\{B|\delta_{s}\subseteq\Delta_{t},|\Delta_{t}|\leq t,|\delta_{s}|\leq s\}|\\ &\leq\sum_{t^{\prime}=0}^{t}\underbrace{|\{B||\Delta_{t}|=t^{\prime}\}|}_{I_{t^{\prime}}}\times\underbrace{|\{B|\delta_{s}\subseteq\Delta_{t},|\Delta_{t}|=t^{\prime},|\delta_{s}|\leq s\}|}_{I_{s}}.\end{split}$ (8)
$I_{t^{\prime}}$ is upper bounded in [17] as
$I_{t^{\prime}}\leq{{2q_{d}+t^{\prime}+1}\choose{2q_{d}+1}}{{q_{d}+t^{\prime}}\choose{q_{d}}}$ (9)
For $I_{s}$, we have
$\begin{split}I_{s}&=|\{B|\delta_{s}\subseteq\Delta_{t},|\Delta_{t}|=t^{\prime},|\delta_{s}|\leq s\}|\\ &=\sum_{s^{\prime}=0}^{s}|\{B|\delta_{s}\subseteq\Delta_{t},|\Delta_{t}|=t^{\prime},|\delta_{s}|=s^{\prime}\}|\\ &\leq\sum_{s^{\prime}=0}^{s}{q_{s}\choose q_{s}-s^{\prime}}{t^{\prime}\choose s^{\prime}},\end{split}$ (10)
which, since (${q_{s}\choose q_{s}-s^{\prime}}\leq{q_{s}+s^{\prime}\choose q_{s}}\leq{q_{s}+s\choose q_{s}}$, and (${t^{\prime}\choose s^{\prime}}\leq{t^{\prime}+s^{\prime}\choose s^{\prime}}\leq{t^{\prime}+s\choose s}$, can be upper bounded as
$\begin{split}I_{s}\leq(s+1){q_{s}+s\choose q_{s}}{t^{\prime}+s\choose s}.\end{split}$ (11)
Summing up, we have
$\begin{split}I\leq&(s+1){q_{s}+s\choose q_{s}}\sum_{t^{\prime}=0}^{t}{{2q_{d}+t^{\prime}+1}\choose{2q_{d}+1}}{{q_{d}+t^{\prime}}\choose{q_{d}}}{t^{\prime}+s\choose s}\\ \leq&(s+1)(t+1){q_{s}+s\choose q_{s}}{t+s\choose s}{{q_{d}+t}\choose{q_{d}}}{{2q_{d}+t+1}\choose{2q_{d}+1}}.\end{split}$ (12)
Definition: For integers $t$ and $s$, we say that $X\in\{0,1\}^{n}$ is $(t,s)$-bad if there exist two patterns $A$ and $B$ such that $|\Delta_{t}|\geq t$ or $|\delta_{s}|\geq s$, and $X_{A}=X_{B}$.
In the following lemma, we provide an upper bound on the number of $(t,s)$-bad sequences.
Lemma 4
Fixing the numbers of deletions and substitutions, the number of $(t,s)$-bad sequences in $X\in\{0,1\}^{n}$, $I_{(t,s)-bad}$, can be upper bounded as
$I_{(t,s)-bad}\leq{n\choose q_{d}}^{2}{n-q_{d}\choose q_{s}}^{2}2^{n-t}.$ (13)
Proof:
We first obtain the probability that a sequence is $(t,s)$-bad, $P_{(t,s)-bad}$. According to Corollary 2, the probability that two patterns $A$ and $B$ lead to the same output is upper bounded by $2^{-|\Delta_{t}|}$. Since this probability is decreasing in $|\Delta_{t}|$, and using union bound trick, we have
$P_{(t,s)-bad}$ $=\!Pr_{X\in U^{\{0,1\}^{n}}}[\exists A,B|(|\Delta_{t}|\!\geq t\lor\!|\delta_{s}|\geq s)\!\land\!X_{A}\!\!=\!\!X_{B}]$
$\leq\sum_{A,B}Pr_{X\in U^{\{0,1\}^{n}}}[X_{A}=X_{B}]$
($\leq{n\choose q_{d}}^{2}{n-q_{d}\choose q_{s}}^{2}2^{-t}$ (14)
Finally, multiplying this probability by the total number of sequences concludes the proof. ∎
Theorem 1
Fixing the numbers of deletions and substitutions, suppose that there exists a decoding algorithm that succeeds with probability at least $\delta\geq 0$. Then, the size of the codebook satisfies
$\log N\leq n-q_{d}-\log{n\choose q_{d}}-\log{n-q_{d}\choose q_{s}}+\log\frac{2}{\delta}+\log\alpha,$ (15)
where
$\alpha=t^{2}\left(2e\right)^{t-1}\left(\frac{5t}{q_{s}}\right)^{q_{s}}\left(\frac{5t}{q_{d}}\right)^{3q_{d}+1},$
with $t=\lceil{3q_{d}}\log\frac{ne}{q_{d}}+{3q_{s}}\log\frac{(n-q_{d})e}{q_{s}}+\log\frac{2}{\delta}\rceil.$
Proof:
We first assume that the receiver knows the numbers of deletions and substitutions as side information. We consider a guesser that, given the received sequence $Z_{A}$, outputs the transmitted sequence $Z$ and the pattern pair $A=(A_{t},A_{s})$, both chosen uniformly at random, with a non-negligible probability. More specifically, on the input $Z_{A}$, the guesser first decodes it with the proposed decoder $R(\cdot)$ with success probability $\delta$, then outputs $g(Z_{A})=(R(Z_{A}),A^{\prime})$, where $A^{\prime}$ is the lexicographically first sequence to satisfy $(R(Z_{A}))_{A^{\prime}}=Z_{A}$ if one exists, and the fixed pattern pair $A^{\prime}=(A_{t}^{\prime}=1,2,\cdots,n-q_{d},A_{s}^{\prime}=1,2,...,q_{s})$ otherwise.
We now lower bound the success probability of the guesser. For this, we first consider the probability that the decoding succeeds ($R(Z_{A})=Z$) and the codeword $Z\in\mathcal{C}$ is not $(t,s)$-bad as follows
[…]
$\geq\delta-\frac{2^{n-t}}{N}{n\choose q_{d}}^{2}{n-q_{d}\choose q_{s}}^{2}$ (16)
since $Pr_{Z\in U^{\mathcal{C}}}[Z\text{ is {(t,s)}-bad}]\leq\frac{|\text{{(t,s)}-badsequences}\in\{0,1\}^{n}|}{|\mathcal{C}|}$.
If $N\leq\frac{2^{n}}{{n\choose q_{d}}{n-q_{d}\choose q_{s}}}$, the theorem follows trivially. For $N\geq\frac{2^{n}}{{n\choose q_{d}}{n-q_{d}\choose q_{s}}}$, we have
$\begin{split}\frac{2^{n-t}}{N}{n\choose q_{d}}^{2}{n-q_{d}\choose q_{s}}^{2}&\leq 2^{-t}{n\choose q_{d}}^{3}{n-q_{d}\choose q_{s}}^{3}\\ &\leq 2^{-t}\left(\frac{ne}{q_{d}}\right)^{3q_{d}}\left(\frac{(n-q_{d})e}{q_{s}}\right)^{3q_{s}},\end{split}$ (17)
making which smaller or equal to $\frac{\delta}{2}$ (our own choice), gives us a lower bound on $t$ as $t\geq{3q_{d}}\log\frac{ne}{q_{d}}+{3q_{s}}\log\frac{e(n-q_{d})}{q_{s}}+\log\frac{2}{\delta}$. We take $t=\lceil{3q_{d}}\log\frac{ne}{q_{d}}+{3q_{s}}\log\frac{(n-q_{d})e}{q_{s}}+\log\frac{2}{\delta}\rceil$.
If $Z$ is not $(t,s)$-bad, pattern $B$ gets mixed up with the desired pattern $A$ only when $\Delta_{t}\leq t-1$ and $\delta_{s}\leq s-1$. According to Lemma 7, the number of such patterns is upper bounded by
$st{q_{s}+s-1\choose q_{s}}{t+s-2\choose s-1}{{q_{d}+t-1}\choose{q_{d}}}{{2q_{d}+t}\choose{2q_{d}+1}}$
$\leq t^{2}{q_{s}+t-1\choose q_{s}}{2t-2\choose t-1}{{q_{d}+t-1}\choose{q_{d}}}{{2q_{d}+t}\choose{2q_{d}+1}}$
$\leq t^{2}\left(e\frac{q_{s}+t-1}{q_{s}}\right)^{q_{s}}\left(2e\right)^{t-1}\left(e\frac{q_{d}+t-1}{q_{d}}\right)^{q_{d}}\left(e\frac{2q_{d}+t}{2q_{d}+1}\right)^{2q_{d}+1}$
$\leq t^{2}\left(2e\right)^{t-1}\left(\frac{5t}{q_{s}}\right)^{q_{s}}\left(\frac{5t}{q_{d}}\right)^{3q_{d}+1}$
$:=\alpha,$ (18)
knowing that $s\leq t$ due to $\delta_{s}\subseteq\Delta_{t}$ for two patterns to be mistaken with each other, and using the fact that (${n\choose k}\leq\left(ne/k\right)^{k}$. We also used $t\geq 3q_{d}$ and $t\geq 3q_{s}$, since $t\geq{3q_{d}}\log\frac{ne}{q_{d}}+{3q_{s}}\log\frac{e(n-q_{d})}{q_{s}}+\log\frac{2}{\delta}$.
Conditioned on the decoding succeeding and the codeword not being $(t,s)$-bad, each deletion/substitution pattern is equally likely, and hence, the lexicographically first pattern is correct with probability at least $\alpha^{-1}$. Hence, the total success probability of the guesser is at least
$Pr_{Z\in U^{\mathcal{C}}}[g(Z_{A})=(Z,A)]\geq\frac{\delta}{2}\alpha^{-1}.$ (19)
Finally, combining this lower bound with the upper bound in Corollary 1, we have
$\frac{\delta}{2}\alpha^{-1}$ $\leq\frac{2^{n-q_{d}}}{N{n\choose q_{d}}{n-q_{d}\choose q_{s}}}$ (20)
which gives us an upper bound on the codebook size for the case with side information at the receiver as
$\log N$ $\leq n-q_{d}-\log{{n\choose q_{d}}}-\log{{n-q_{d}\choose q_{s}}}+\log\frac{2}{\delta}+\log\alpha.$ (21)
Since side information cannot decrease the capacity, this is also an upper bound on the codebook size of the case without side information. This concludes the proof. ∎
IV Random numbers of deletions and substitutions
Going from the case of fixed numbers of deletions and substitutions to the case where the numbers of deletions and substitutions are random merely involves taking advantage of the concentration of the numbers of deletions and substitutions around their means, $np_{d}$ and $n(1-p_{d})p_{s}^{\prime}$, respectively.
Theorem 2
In the setup with random numbers of deletions and substitutions, suppose that there exists a decoding algorithm that succeeds with probability at least $\delta$ for codeword length of $n\geq\frac{12\log\frac{4}{\delta}}{\min\{p_{d},(1-p_{d})p_{s}^{\prime}\}}$. Then, the codebook size $N$ satisfies
$\begin{split}\log N\leq n-q_{d,\min}-\log{n\choose q_{d,\min}}&-\log{n-q_{d,\max}\choose q_{s,\min}}\\ &+\log\frac{8}{\delta^{2}}+\log\alpha_{\max}^{*},\end{split}$ (22)
where $q_{d,\min}=(1-\gamma)np_{d}$, $q_{s,\min}=(1-\gamma)n(1-p_{d})p_{s}^{\prime}$, $q_{d,\max}=(1+\gamma)np_{d}$, $q_{s,\max}=(1+\gamma)n(1-p_{d})p_{s}^{\prime}$, with Let $\gamma=\sqrt{\frac{3\log\frac{4}{\delta}}{np_{d}}}$, and
$\alpha_{\max}^{*}={t_{\max}^{*2}}\left(2e\right)^{t_{\max}^{*}-1}\left(\frac{5t_{\max}^{*}}{q_{s,\max}^{*}}\right)^{q_{s,\max}^{*}}\left(\frac{5t_{\max}^{*}}{q_{d,\max}^{*}}\right)^{3q_{d,\max}^{*}+1},$ (23)
with $t_{\max}^{*}=\lceil{3q_{d,\max}^{*}}\log\frac{ne}{q_{d,\max}^{*}}+{3q_{s,\max}^{*}}\log\frac{e(n-q_{d,\max}^{*})}{q_{s,\max}^{*}}+\log\frac{8}{\delta^{2}}\rceil$.
Proof:
Suppose that we have a decoding algorithm for the deletion/substitution channel that succeeds on codebook $\mathcal{C}$ with probability $\delta\geq 0$. The standard multiplicative Chernoff bound gives an upper bound on the probability that a random variable lies outside an interval around its mean. More specifically (such as [19][Corollary 4.6]), if $X_{i}$’s are Bernoulli random variables with success probability $p_{i}$, then, $Pr(|\sum_{i}X_{i}-\mu|\geq\gamma\mu)\leq 2e^{-\frac{\mu\gamma^{2}}{3}}$, where $\mu=\sum_{i}p_{i}$. With $q_{d}=\sum_{i}X_{i}$, which is the random variable representing the number of deletions, we have $Pr(|q_{d}-np_{d}|\leq\gamma np_{d})\geq 1-2e^{-\frac{np_{d}\gamma^{2}}{3}}$. Setting the probability that the number of deletions, with mean $\mu_{d}=np_{d}$, lies in the interval $(1-\gamma)np_{d}\leq q_{d}\leq(1+\gamma)np_{d}$ to be at least $1-\frac{\delta}{2}$, we have $1-2e^{-\frac{np_{d}\gamma^{2}}{3}}=1-\frac{\delta}{2}$. This gives us $\gamma=\sqrt{\frac{3\log\frac{4}{\delta}}{np_{d}}}$. Finally, setting $\gamma\leq\frac{1}{2}$, as a design choice, we get, $n\geq\frac{12\log\frac{4}{\delta}}{np_{d}}$. Similarly, for the probability of the number of substitutions, with mean $\mu_{s}=n(1-p_{d})p_{s}^{\prime}$, lying in the interval $(1-\gamma)n(1-p_{d})p_{s}^{\prime}\leq q_{s}\leq(1+\gamma)n(1-p_{d})p_{s}^{\prime}$ to be at least $1-\frac{\delta}{2}$, we need $n\geq\frac{12\log\frac{4}{\delta}}{(1-p_{d})p_{s}^{\prime}}$. Combining the two conditions, we finally have $n\geq\frac{12\log\frac{4}{\delta}}{\min\{p_{d},(1-p_{d})p_{s}^{\prime}\}}$, as stated in the theorem statement.
Next, since deletions and substitutions are independent, we have
$Pr\{(\text{Successful decoding})$
$\hskip 28.45274pt\land(|q_{d}-\mu_{d}|\leq\gamma\mu_{d}\cap|q_{s}-\mu_{s}|\leq\gamma\mu_{s})\}$
$\geq Pr\{\text{Successful decoding}\}$
$\hskip 28.45274pt+Pr\{|q_{d}-\mu_{d}|\leq\gamma\mu_{d}\cap|q_{s}-\mu_{s}|\leq\gamma\mu_{s}\}-1$
$=Pr\{\text{Successful decoding}\}$
$\hskip 28.45274pt+Pr\{|q_{d}-\mu_{d}|\leq\gamma\mu_{d}\}Pr\{|q_{s}-\mu_{s}|\leq\gamma\mu_{s}\}-1$
$\geq\frac{\delta^{2}}{4}.$ (24)
Hence, there must be some $(q_{d}^{*},q_{s}^{*})$ in the defined intervals such that the success probability of the exact $(q_{d}^{*},q_{s}^{*})$ deletion/substitution channel is at least $\frac{\delta^{2}}{4}$.
Let $\alpha^{*}$ be given by
$\alpha^{*}={t^{*}}^{2}\left(2e\right)^{t^{*}-1}\left(\frac{5t^{*}}{q_{s}^{*}}\right)^{q_{s}^{*}}\left(\frac{5t^{*}}{q_{d}^{*}}\right)^{3q_{d}^{*}+1},$ (25)
where
$t^{*}=\lceil{3q_{d}^{*}}\log\frac{ne}{q_{d}^{*}}+{3q_{s}^{*}}\log\frac{e(n-q_{d}^{*})}{q_{s}^{*}}+\log\frac{8}{\delta^{2}}\rceil.$ (26)
Note that $\alpha^{*}$ is a non-decreasing function of both $q_{d}$ and $q_{s}$, hence it is maximized with their maximum values. Finally, using the result in Theorem 1, by putting in each term the values in the intervals that maximize it, concludes the proof. ∎
Theorem 3
The capacity of a deletion/substitution channel can be upper-bounded as
$C\leq 1-(1-o(1))(H(p_{d})+H(p_{s}^{\prime})).$ (27)
Proof:
Consider any fixed $p_{d}$ with $n$ going to infinity. We have $-n(1-\gamma)p_{d}\log((1-\gamma)p_{d})\leq\log{n\choose(1-\gamma)np_{d}}\leq-n(1-\gamma)p_{d}\log(\frac{1-\gamma}{e}p_{d})$. So $\log{n\choose(1-\gamma)np_{d}}=n(H(p_{d})+o(1))$, as $\gamma$ is vanishing. Similarly, for a any fixed $p_{s}$, as $p_{d}$ goes to zero, we have $p_{s}^{\prime}\rightarrow p_{s}$. So, we have $\log{n-(1+\gamma)np_{d}\choose(1-\gamma)n(1-p_{d})p_{s}^{\prime}}=n(H(p_{s})+o(1))$. Further, it can be shown that $\log\alpha_{\max}^{*}=\mathcal{O}(n(p_{d}+p_{d})\log(\log(1/p_{d})+\log(1/p_{s}^{\prime})))=o(n(H(p_{d})+H(p_{s}))$ as $p_{d}$ and $p_{d}$ (hence, also $p_{s}^{\prime}$) go to 00. Therefore, the three terms $n-\log{n\choose(1-\gamma)np_{d}}-\log{n-(1+\gamma)np_{d}\choose(1-\gamma)n(1-p_{d})p_{s}^{\prime}}$ dominate the right-hand side of the equation; Dividing both sides by $n$, we obtain $\frac{1}{n}\log N\leq 1-(1-o(1))(H(p_{d})+H(p_{s}))$, which completes the proof. Note that the $o(1)$ term is understood as going to 0 in the limit as first $n$ goes to infinity and then as $p_{d}$ and $p_{s}$ go to 00. ∎
Corollary 3
Combined with the lower bound result in [19], it can be observed that the capacity of the deletion/substitution channel approaches $1-H(p_{d})-H(p_{s})$ for the probabilities of deletion $p_{d}$ and substitution $p_{s}$ near zero.
V Conclusion
In this paper, we characterized the capacity of binary channels with synchronization errors, where transmitted bits are randomly and independently deleted and flipped. To this end, we first obtained an upper bound on the capacity, assuming that the numbers of deletions and substitutions are known at the receiver. We then extended the results to the case of random numbers of deletions and substitutions. Finally, combining with an existing lower bound, we showed that, in the regime of asymptotically small deletion and substitution probabilities, the rate loss is equal to the sum of the binary entropies of the deletion and substitution probabilities. As a future work, it would be of interest to extend this result to channels with insertion, deletion, and substitution (IDS) errors.
Acknowledgment
The work of Mohammad Kazemi was supported by UKRI under the U.K. government’s Horizon Europe Funding Guarantee under Grant 101103430. Tolga M. Duman’s work was funded by the European Union ERC TRANCIDS 101054904. Views and opinions expressed are, however, those of the author(s) only and do not necessarily reflect those of the European Union or the European Research Council Executive Agency. Neither the European Union nor the granting authority can be held responsible for them.